OpenStudy (anonymous):

1. Given f(x) = 3x^6, find f ^-1(x). Then, state whether f ^-1(x) is a function.

OpenStudy (dape):

Start with substituting \(f(x)\) for x and x for y, then solve for y, the y that you get is your inverse \(f^{-1}(x)\).

OpenStudy (dape):

So the equation you get after the substitutions is \[ x=3y^6 \] Solve this for y.

OpenStudy (anonymous):

is it a function?

OpenStudy (dape):

Have you solved it?

OpenStudy (anonymous):

y = +- (x/3) ^6.

OpenStudy (dape):

That's almost right, you just get the exponent 1/6, so \[ y=f^{-1}(x)=\pm\left(\frac{x}{3}\right)^{1/6} \] This is not really a function since for a single value of x, you always get 2 values for the inverse (one for plus and one for minus). One can make it a function though, by ignoring either the positive or the negative value.

OpenStudy (dape):

So you could say it's a function if you look at positive x's.

OpenStudy (anonymous):

Thanks a lot